A block of mass m = 2.00 kg is released from rest at the top of an inclined plane. The block starts out at height h = 0.100 m above the top of the table, the table height is H = 2.00 m, and θ = 40.0°.

Find the acceleration of the block while it slides down the incline.
1
Expert's answer
2018-10-01T13:46:09-0400
F=ma
F=mg·sin⁡α
ma=mg·sin⁡α
a=g·sin⁡α=9.81·sin⁡〖40 ̊=9.81·0.643=6.3 m/s^2 〗